With the help of sequences, calculate √5 correct upto 5 decimals.
1
Expert's answer
2020-04-27T18:32:27-0400

"\sqrt 5=\sqrt{x*(5/x)}"

Now using the AM-GM inequality, we get;

"(x+(5/x))/2 \ge \sqrt{x*(5/x)}"

"\implies \sqrt 5 \le (x+(5/x))/2"

Thus, using this inequality as an approximation we can correlate the following iterative formula for calculating the value of "\sqrt 5" as follows :


Let us consider a recursive sequence given by the formula : "x_{n+1}=(x_n+5/x_n)/2 ; \forall n \ge2; x_1=2"

"\implies x_2=(2+5/2)/2=2.25"

"\implies x_3=(2.25+5/2.25)/2=2.236111"

"\implies x_4=(2.236111+5/2.236111)/2=2.236068"

"\implies x_5=(2.236068+5/2.236068)/2=2.236068"


Thus, the sequence converges to the value "2.236068" .

Exact value : "\sqrt 5=2.2360679775..." "\approx 2.23606" (correct upto 5 decimal places)

"x_4=2.236068 \approx 2.23606" (upto 5 decimal places)


Thus, we obtain the required value upto the required precision in 3 iterations only using the above defined sequence.